a spire in the form of a regular heptagonal pyramid whose base edge is 8 ft. and whose altitude is 75 ft. is to be at a cost of 22 cents per square yard. what is the total cost?
Answer by Alan3354(69443)

a spire in the form of a regular heptagonal pyramid whose base edge is 8 ft. and whose altitude is 75 ft. is to be at a cost of 22 cents per square yard. what is the total cost?
-----------
Find the lateral area.
---
Each face is a triangle.
The slant height is the hypotenuse of a right triangle with legs of 75 feet and the apothem of the heptagon.
Find the apothem a:
a = (s/2)/tan(360/14)
a =~ 8.306
------
SH = sqrt(a^2 + 75^2)
SH =~ 75.4585 ft.
-----
Area of each triangle = b*h/2
= 8*SH/2
= 301.834
----
LA = 7*301.834
= 2112.839 sq ft
= 234.76 sq yards
---
Cost = LA*$0.22
= $51.65